The Last problem for the summer
25^(x+2)=5^(3x-4)
I think you take like 5 out and make 5 5
so then its 5x+2=3x-4 then maybe 2x=-2 then divide by two so x is negative one but i think its wrong any help guys?

Answer by stanbon(75887)

25^(x+2)=5^(3x-4)
(5^2)^(x+2) = 5^(3x-4
2x+4 = 3x-4
x = 8
===========
Cheers,
Stan H.